OpenStudy (anonymous):

MEDAL FOR THE BEST ANSWER Which is an irrational number? A. / -3 / B. 0.6 ____ C. V 0/9 (square root 0/9 ) ______ D. V 5 /3 (square root 5/3 )

OpenStudy (zzr0ck3r):

anything that you cant write as the ratio of two integers.

OpenStudy (zzr0ck3r):

|-3| = 3 = 3/1 so nope .6 = 6/10 so nope sqrt(0/9) = 0 = 0/1 so nope sqrt{5/3} = sqrt{5}\sqrt{3} ding ding ding

OpenStudy (zzr0ck3r):

square root of a prime number is always irrational.

OpenStudy (anonymous):

Thank you!!